OpenStudy (anonymous):

Write the quadratic function f(x) = x^2 + 8x + 3 in vertex form. A)f(x) = (x - 4)^2 - 13 B)f(x) = (x - 4)^2 + 3 C)f(x) = (x + 4)^2 + 3 D)f(x) = (x + 4)^2 - 13

OpenStudy (anonymous):

when you completed the square you didn't subtract 16 should've been x^2+8x+16+3-16

OpenStudy (anonymous):

so the answer would be D

OpenStudy (anonymous):

f'(x) = 2x + 8 2x + 8 = 0 x = -4 f(-4) = (-4)^2+8(-4) +3 f(-4) = 16 + -32 + 3 = -13 Vertex is (-4,-13) So answer should be D.

OpenStudy (anonymous):

moneybird you could have just completed the square instead of using calculus lol

OpenStudy (anonymous):

whatever, complete the square is long...